# Designing and Evaluating AI Margin Notes in Document Reader Software

### v1

@@ Abstract.
AI capabilities for document reader software are usually presented in separate chat interfaces. We explore integrating AI into document comments, a concept we formalize as AI margin notes. Three design parameters characterize this approach: margin notes are integrated with the text while chat interfaces are not; selecting text for a margin note can be automated through AI or manual; and the generation of a margin note can involve AI to various degrees. Two experiments investigate integration and selection automation, with results showing participants prefer integrated AI margin notes and manual selection. A third experiment explores human and AI involvement through six alternative techniques. Techniques with less AI involvement resulted in more psychological ownership, but faster and less effortful designs are generally preferred. Surprisingly, the degree of AI involvement had no measurable effect on reading comprehension. Our work shows that AI margin notes are desirable and contributes implications for their design.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; copyright: noneccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction techniques
@@
[…]
Side-by-side comparison of two interfaces for describing text about flowering plants. Left: AI margin note highlights text, providing an explanation beside the document. Right: Chat-based interface with a chatbot explaining the same concept. Both focus on floral structures with a modern, educational tone.
@@ 1. Introduction
Taking notes while reading documents is a common active reading strategy that can be done in two primary ways. First, people can write notes that are decoupled from the document text by writing on a separate piece of paper. Second, people can write notes that are integrated with the document text (Adler et al., 1998), for example, by writing in the margins (marginalia). Such ‘margin notes’ have been common practice for centuries, as they provide a space for readers to summarize, paraphrase, explain unfamiliar concepts, make connections to existing knowledge, and even express personal opinions (Jackson, 2001; Porter-O’Donnell, 2004). This can be especially beneficial when reading for educational purposes (Bold and Wagstaff, 2017), as it provides space for people to work through difficult sections and quickly re-access their thoughts afterwards within a shared context (Marshall, 1997; Adler et al., 1998). When reading digital documents, margin notes can be created using pen input devices (e.g., (Romat et al., 2019)), but a more common way is by leaving comments in systems like Google Docs, Microsoft Word, and Adobe Acrobat Reader (Wolfe and Neuwirth, 2001). This is usually done by selecting text within a document and writing in a text box positioned beside it.
More recently, people have been leveraging the capabilities of large language models (LLMs) to support reading documents, for example, to summarize, reword, and better understand unfamiliar concepts (Guidroz et al., 2025; Gu et al., 2024; Kreijkes et al., 2025). The generated output serves a purpose like margin notes, yet the capabilities of LLMs have not been integrated into commenting features of document reader software. Instead, most LLM-enhanced document readers require users to summarize and ask questions about the text in a separate, chat-like interface disconnected from the document. There is an opportunity to improve how LLMs are used in this context by integrating their capabilities directly into ‘margin note’ comments that are linked to the document text.
@@
Our results indicated that AI margin notes were preferred over chat interfaces and that selecting text for AI margin notes should be done manually. AI margin note techniques with varying levels of human and AI involvement were valued for different reasons. For example, techniques with more human involvement were typically associated with more psychological ownership, but techniques with more AI involvement were faster, less effortful, and generally preferred. Our work contributes the idea of AI margin notes and empirical results demonstrating that AI margin notes are a desirable feature for document reader software.
@@ 2.1. Note-Taking while Reading
There are many types of notes that readers can create, margin notes being one of them. Research in psychology suggests that the general activity of note-taking is beneficial for two main reasons (Kiewra, 1989). First, notes act as external storage for information, which can be re-read to reinforce memory (the storage function). Second, the act of creating notes can facilitate learning as it requires readers to pay more attention to the material to process and organize ideas (the encoding function).
The encoding function is especially beneficial when the reader processes the text at a deeper level (Kiewra, 1985; Craik and Lockhart, 1972), for example, by connecting it to prior knowledge and experiences. However, many people opt for shallower and less effective note-taking strategies. For example, Bretzing and Kulhavy’s analysis of students’ note-taking activities (Bretzing and Kulhavy, 1981) revealed that students tend to take verbatim notes that repeats text from documents. In another study (Bretzing and Kulhavy, 1979), they showed that students who took verbatim notes performed worse on a test than those who used deeper note-taking strategies by writing their own summaries or by paraphrasing text from the document.
Note-taking is a complex activity that requires significant cognitive effort as readers must coordinate and frequently switch between reading and writing activities. Writing takes more time than reading, and excessive delays between reading activities as a result of note-taking can hinder comprehension. Therefore, readers often experience significant mental and temporal demand, even when they are reading and taking notes without any time limits (Piolat et al., 2005). This can become especially tiring with longer (Kobayashi, 2005) or poorly-formatted documents (Olive and Barbier, 2017). Although researchers have argued that increased cognitive effort can benefit learning (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a) and memory (Tyler et al., 1979), it does not always lead to improved learning outcomes, for example, if the task is too frustrating or if the learner lacks motivation (Grund et al., 2024).
@@ 2.2. Reading and Note-Taking with LLMs
Recent work has investigated how LLMs can improve user experience and comprehension while reading and taking notes. For example, text simplification techniques that turn complex documents into simplified versions (Guidroz et al., 2025) and make documents easier to skim (Gu et al., 2024) can improve reading comprehension and reduce workload. Users of systems like ChatGPT, Adobe Acrobat (Adobe, 2025), Google Notebook LM (Google, 2025), NoteGPT (NoteGPT, 2025), and ChatPDF (ChatPDF, 2023) can ask questions, summarize, and write notes about documents. However, few investigations examine the effect on factors like reading comprehension, reading duration, workload, and preferences.
@@
Kreijkes et al. (Kreijkes et al., 2025) asked high school students to try different note-taking techniques to study for reading comprehension tests that took place three days later: note-taking independently and note-taking while having access to an LLM-powered chatbot to ask questions. Both of these note-taking techniques were compared to a baseline of asking the chatbot questions about the document, without any note-taking. Their results showed that both note-taking techniques led to better performance than just asking the chatbot questions about the document, suggesting that using LLMs in a more cognitively engaging way (i.e., with some note-taking) can improve reading comprehension. However, they did not compare the two note-taking conditions, making the effect of LLM use on note-taking unclear.
@@
Although Kreijkes et al. found that just asking the LLM questions led to poorer test performance, it was preferred by participants as it was perceived to be more enjoyable and less effortful. Such findings have been reproduced in other studies focused on the effects of LLMs on learning more broadly. Systems like ChatGPT can allow for personalized learning experiences, which can improve academic performance, reduce mental effort, and motivate learners (Wang and Fan, 2025; Deng et al., 2025), however, learners may over-rely on these systems (Wang and Fan, 2025). A balance may be to encourage more cognitive engagement. For example, ChatPRCS (Wang et al., 2024) generates practice reading comprehension questions for students, which increased mental load but improved reading comprehension. Similarly, CoAsker (Liu et al., 2023) encourages students to generate practice questions with an LLM-powered assistant, which were displayed in a side panel much like notes beside the document. When compared to generating practice questions without assistance, receiving questions that were generated by the assistant led to higher reading comprehension scores.
@@
To our knowledge, no prior work has thoroughly explored integrating LLMs into the commenting feature of document reader software. The closest prior work are prototype mockups in Melin-Higgins’ bachelor’s thesis (Melin-Higgins, 2024), in which users select document text to issue pre-determined prompts to a hypothetical LLM-powered assistant. Responses appear as ‘sticky notes’ beside the selected text, like margin notes, or they appear underneath the selected text by modifying the document structure. A very small 3-person study showed a preference for the margin note style. The study also only focused on usability and user preferences, and did not consider factors like reading comprehension. Furthermore, the prototypes were partially functioning Figma mockups, and none explored different levels of human and AI involvement.
@@ 3.1. Integration
The primary difference between AI margin notes and chat-based tools is integration. Specifically, AI margin notes enhance comments in document readers. Comments are anchored to specific text, making them integrated into the document, which contrasts with chat-based tools that are placed in a separate side panel with the content disconnected from the context (Figure 2). A separate interface for interacting with an AI assistant may suffice for general questions about the document or to receive overall summaries. However, readers also direct the AI assistant to specific parts of the text to contextualize their prompt (Kreijkes et al., 2025). Separating these responses from the text can be inefficient.
First, referring to specific parts of a document while prompting requires additional work (Masson et al., 2024). For example, if the user wants to simplify the language of a paragraph in a scientific document (e.g., (Guidroz et al., 2025)), they must first type a command like “simplify this,” select the paragraph from the document, then copy and paste it after the command. This copy and paste step could be avoided entirely if the user selected the paragraph and typed “simplify this” in an AI margin note. Or easier, by just selecting text and pressing a button that reifies the “simplify this” prompt (Masson et al., 2024; Beaudouin-Lafon and Mackay, 2000).
@@
Second, shifting attention to a separate interface may distract from reading. Note-taking requires frequent switching between reading and writing (Adobe, 2025; Piolat et al., 2005), and researchers suggest that note-taking should “interrupt reading as little as possible” (Marshall, 1997). Having readers frequently switch between separate reading and writing interfaces may further increase cognitive load (Harrison et al., 1995). These switching costs could be mitigated by presenting the prompting interface alongside the specific text the user is reading.
@@
Third, referring back to specific responses requires additional work. For example, suppose a user used a chat-based LLM interface to get an explanation for an unfamiliar concept in a document. The explanation may be remembered in the short term, but if the user revisits the document much later, they must scroll through a lengthy and poorly organized chat history to find it, which may be especially difficult for readers with a lower working memory capacity to do (Sanchez and Wiley, 2009). Linking responses directly to the relevant text avoids this issue. Some systems, like Adobe Acrobat Reader, support clicking on chat responses to highlight relevant parts of the document, however, this still requires scrolling through chat history. Alternatively, the user could re-issue the prompt, but this wastes time and resources. Of course, users could copy responses they wish to save and paste them in comments (Kreijkes et al., 2025), but this requires additional copy and paste steps that would not be needed with an AI margin note.
Figure 2. Integration: (a) a chat-based interface is not integrated since it is separated from the associated document text and (b) AI margin notes are integrated since they are associated to specific document text. In this and following two figures: green denotes human-written text, like a prompt, and blue represents AI-generated text, like a response. Yellow indicates text that the AI margin note is linked to.
[…]
Conceptual diagram comparing two layouts. (a) Separated: text on left, colored boxes on right. (b) Integrated: text and boxes combined, aligned side by side.
@@ 3.2. Selection Automation
Figure 3. Selection automation: (a) text can be automatically selected and multiple AI margin notes created at once and (b) the user can manually select text to create an AI margin note.
[…]
Conceptual diagram comparing two methods: (a) Automated summarization with a ”Summarize” button, arrow to a highlighted text box. (b) Manual with text selection, ”Add Summary” button, and similar output.
Figure 4. Levels of human and AI involvement: traditional margin notes consist of text that is 100% written by a human, but an AI margin note could consist of text that was 100% generated by AI, or could involve more human-generated text.
[…]
Conceptual diagram showing a shift from 100% AI with blue notes to 100% human with green notes, labeled ”AI margin notes with more human involvement.”
AI margin notes are associated with specific text, which defines the context for the prompt and a location to display the note. Specifying text for the note could be done automatically by the AI assistant or manually by the user (Figure 3). For example, consider summarization, a common task readers do in conventional margin notes (Jackson, 2001; Porter-O’Donnell, 2004) and with LLM systems (Melin-Higgins, 2024; Kreijkes et al., 2025). When using LLMs in document reader software, readers can generate a summary of the entire document automatically by pressing a single button (Adobe, 2025), or they can manually specify which parts of the document need to be summarized by copying and pasting text into the chat. When the reader just presses a button, the LLM decides which information is important and worth including in the summary, but when the reader copies text to summarize, they are deciding which parts of the document are most important. Prior work on text highlighting suggests that this decision process could improve learning outcomes and reading comprehension, but it could also increase mental effort (Yue et al., 2015; Joshi and Vogel, 2024).
Automating the selection of text to create AI margin notes may also impact psychological ownership, feelings of the learning experiences belonging to them (Pierce et al., 2001). Prior work shows that fostering psychological ownership, which can be achieved by giving learners more control over their learning (Pierce et al., 2001, 2003), can encourage more active learning (Griffith, 1990) and motivates learning (Soslau and Yost, 2007). Automatically generating AI margin notes reduces effort, but potentially at the risk of lowering comprehension and psychological ownership.
@@ 3.3. Human and AI Involvement
Currently, commenting in document reader software is intended for text that is 100% written by a human. Specifically, the user must formulate their own ideas and type into a text box. At the other extreme, pressing a “Summarize” button in a document reader (Adobe, 2025) produces text that is 100% written by the LLM and does not require any guidance from the human. However, there are techniques for producing text that require involvement from both the human and the LLM. For example, chat-based interfaces require the user to write a prompt so the LLM can produce relevant output text. Here, the user exerts control over what the output should be, and the LLM produces it (Joshi and Vogel, 2025b). These roles can be reversed too, for example, the LLM could ‘prompt’ the user to respond to practice reading comprehension questions (Wang et al., 2024; Liu et al., 2023).
Based on these existing variations of human and AI involvement when interacting with LLMs, we believe that AI margin notes could also vary in how much human and AI involvement they require to produce the final comment text (Figure 4). Techniques that require more human involvement likely require more cognitive engagement than those with more AI involvement, but this may be beneficial for reading comprehension (Bjork, 1994b; Bjork and Bjork, 2011; Bjork, 1994a; Tyler et al., 1979), may discourage readers from taking verbatim notes (Bretzing and Kulhavy, 1979), and may even improve feelings of psychological ownership (Joshi and Vogel, 2025b, a). However, frustrating cognitive engagement can discourage learners (Grund et al., 2024), which may be mitigated by increasing AI involvement.

### v3

@@ Abstract.
AI capabilities for document reader software are usually presented in separate chat interfaces. We explore integrating AI into document comments, a concept we formalize as AI margin notes. Three design parameters characterize this approach: margin notes are integrated with the text while chat interfaces are not; selecting text for a margin note can be automated through AI or manual; and the generation of a margin note can involve AI to various degrees. Two experiments investigate integration and selection automation, with results showing participants prefer integrated AI margin notes and manual selection. A third experiment explores human and AI involvement through six alternative techniques. Techniques with less AI involvement resulted in more psychological ownership, but faster and less effortful designs were generally preferred. Surprisingly, the degree of AI involvement had no measurable effect on reading comprehension. Our work shows that AI margin notes are desirable and contributes implications for their design.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790786isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interaction techniques
@@ 1. Introduction
Taking notes while reading documents is a common active reading strategy that can be done in two primary ways. First, people can write notes that are decoupled from the document text by writing on a separate piece of paper. Second, people can write notes that are integrated with the document text (AdlerDiaryStudy), for example, by writing in the margins (marginalia). Such ‘margin notes’ have been common practice for centuries, as they provide a space for readers to summarize, paraphrase, explain unfamiliar concepts, make connections to existing knowledge, and even express personal opinions (jackson2001marginalia; porter2004beyond). This can be especially beneficial when reading for educational purposes (bold2017marginalia), as it provides space for people to work through difficult sections and quickly re-access their thoughts afterwards within a shared context (marshall1997annotation; AdlerDiaryStudy). With digital documents, margin notes can be added as digital ink with a pen (e.g., (romat2019spaceink)), but a more common way is leaving comments when reading documents in systems like Google Docs, Microsoft Word, and Adobe Acrobat Reader (wolfe2001margins). This is usually done by selecting text within a document and writing in a text box positioned beside it.
Large language models (LLMs) can support reading documents to summarize, reword, and better understand unfamiliar concepts (GoogleGuidroz2025llm; GuTextSmimming2024; kreijkes2025effects). The generated output serves a purpose like margin notes, yet the capabilities of LLMs have not been integrated into commenting features of document reader software. Instead, most LLM-enhanced document readers use a separate, chat-like interface disconnected from the document. There is an opportunity to improve how LLMs are used in this context by integrating their capabilities directly into ‘margin note’ comments that are linked to the document text.
@@
Our results indicated that AI margin notes were preferred over chat interfaces and that selecting text for AI margin notes should be done manually. AI margin note techniques with varying levels of human and AI involvement were valued for different reasons. For example, techniques with more human involvement were typically associated with more psychological ownership, but techniques with more AI involvement were faster, less effortful, and generally preferred. Our work contributes:
  • the idea of AI margin notes: comments that are enhanced with LLMs to support note-taking that is integrated with the document text; and
  • empirical results from three experiments, each focusing on a specific design parameter, demonstrating that AI margin notes are a desirable feature for document reader software.
@@ 2.1. Note-Taking while Reading
There are many types of notes that readers can create, margin notes being one of them. Research in psychology suggests that the general activity of note-taking is beneficial for two main reasons (KiewraNoteTakingReview). First, notes act as external storage for information, which can be re-read to reinforce memory (the storage function). Second, the act of creating notes can facilitate learning as it requires readers to pay more attention to the material to process and organize ideas (the encoding function).
The encoding function is especially beneficial when the reader processes the text at a deeper level (kiewra1985investigating; craik1972levels), for example, by connecting it to prior knowledge and experiences. However, many people opt for shallower and less effective note-taking strategies. For example, Bretzing and Kulhavy’s analysis of students’ note-taking activities (bretzing1981note) revealed that students tend to take verbatim notes that repeats text from documents. In another study (bretzing1979processing), they showed that students who took verbatim notes performed worse on a test than those who used deeper note-taking strategies by writing their own summaries or by paraphrasing text from the document.
Note-taking is a complex activity that requires significant cognitive effort as readers must coordinate and frequently switch between reading and writing activities. Writing takes more time than reading, and excessive delays between reading activities as a result of note-taking can hinder comprehension. Therefore, readers often experience significant mental and temporal demand, even when they are reading and taking notes without any time limits (piolat2005cognitive). This can become especially tiring with longer (kobayashi2005limits) or poorly-formatted documents (olive2017processing). Although researchers have argued that increased cognitive effort can benefit learning (bjork1994memory; bjork2011making; bjork1994institutional) and memory (tyler1979cognitive), it does not always lead to improved learning outcomes, for example, if the task is too frustrating or if the learner lacks motivation (grund2024learning).
@@ 2.2. Reading and Note-Taking with LLMs
Recent work has investigated how LLMs can improve user experience and comprehension while reading and taking notes. For example, text simplification techniques that turn complex documents into simplified versions (GoogleGuidroz2025llm; AugustPaperPlain2023) and make documents easier to skim (GuTextSmimming2024) can improve reading comprehension and reduce workload. Users of systems like ChatGPT, Adobe Acrobat (AdobeAcrobatAI), Google Notebook LM (notebooklmGoogleNotebookLM), NoteGPT (NoteGPT), and ChatPDF (chatpdfChatPDFChat) can ask questions, summarize, and write notes about documents. However, few investigations examine the effect on factors like reading comprehension, reading duration, workload, and preferences.
@@
Kreijkes et al. (kreijkes2025effects) asked high school students to try different note-taking techniques to study for reading comprehension tests that took place three days later: note-taking independently and note-taking while having access to an LLM-powered chatbot to ask questions. Both of these note-taking techniques were compared to a baseline of asking the chatbot questions about the document, without any note-taking. Their results showed that both note-taking techniques led to better performance than just asking the chatbot questions about the document, suggesting that using LLMs in a more cognitively engaging way (i.e., with some note-taking) can improve reading comprehension. However, they did not compare the two note-taking conditions, making the effect of LLM use on note-taking unclear.
@@
Although Kreijkes et al. found that just asking the LLM questions led to poorer test performance, it was preferred by participants as it was perceived to be more enjoyable and less effortful. Such findings have been reproduced in other studies focused on the effects of LLMs on learning more broadly. Systems like ChatGPT can allow for personalized learning experiences, which can improve academic performance, reduce mental effort, and motivate learners (wang2025effect; deng2025chatGPT), however, learners may over-rely on these systems (wang2025effect). A balance may be to encourage more cognitive engagement. For example, ChatPRCS (wang2024chatprcs) generates practice reading comprehension questions for students, which increased mental load but improved reading comprehension. Similarly, CoAsker (liu2023student) encourages students to generate practice questions with an LLM-powered assistant, which were displayed in a side panel much like notes beside the document. When compared to generating practice questions without assistance, receiving questions that were generated by the assistant led to higher reading comprehension scores.
@@
To our knowledge, no prior work has thoroughly explored integrating LLMs into the commenting feature of document reader software. The closest prior work are prototype mockups in Melin-Higgins’ bachelor’s thesis (melin2024ai), in which users select document text to issue pre-determined prompts to a hypothetical LLM-powered assistant. Responses appear as ‘sticky notes’ beside the selected text, like margin notes, or they appear underneath the selected text by modifying the document structure. A very small 3-person study showed a preference for the margin note style. The study also only focused on usability and user preferences, and did not consider factors like reading comprehension. Furthermore, the prototypes were partially functioning Figma mockups, and none explored different levels of human and AI involvement.
@@
Figure 2. Integration: (a) a chat-based interface is not integrated since it is separated from the associated document text and (b) AI margin notes are integrated since they are associated to specific document text. In this and the following two figures: green denotes human-written text, like a prompt, and blue represents AI-generated text, like a response. Yellow indicates text that the AI margin note is linked to.
@@ 3.1. Integration
The primary difference between AI margin notes and chat-based tools is integration. Specifically, AI margin notes enhance comments in document readers. Comments are anchored to specific text, making them integrated into the document, which contrasts with chat-based tools that are placed in a separate side panel with the content disconnected from the context (Figure 2). A separate interface for interacting with an AI assistant may suffice for general questions about the document or to receive overall summaries. However, readers also direct the AI assistant to specific parts of the text to contextualize their prompt (kreijkes2025effects). Separating these responses from the text can be inefficient.
First, referring to specific parts of a document while prompting requires additional work (DirectGPT). Consider prompting an LLM to simplify a paragraph in a scientific document (e.g., (GoogleGuidroz2025llm; AugustPaperPlain2023)). With a chat-based interface, the user must carefully formulate a prompt to refer to the specific paragraph (e.g., “simplify the third paragraph in section 3”). Or, the user must select, copy, and paste the paragraph into the chat and write a prompt that uses deictic words to refer to the paragraph instead (e.g., “simplify this”). With an AI margin note, the interaction is simpler as user can use deictic words without a separate copy and paste stage: they just select the paragraph and type “simplify this” where it appears in the document.
@@
Second, shifting attention to a separate interface may distract from reading. Note-taking requires frequent switching between reading and writing (AdobeAcrobatAI; piolat2005cognitive), and researchers suggest that note-taking should “interrupt reading as little as possible” (marshall1997annotation). Having readers frequently switch between separate reading and writing interfaces may further increase cognitive load (harrison1995transparent). These switching costs could be mitigated by presenting the prompting interface alongside the specific text the user is reading.
Figure 3. Selection automation: (a) text can be automatically selected and multiple AI margin notes can be created at once and (b) the user can manually select text to create an AI margin note.
Figure 4. Levels of human and AI involvement: traditional margin notes consist of text that is 100% written by a human, but an AI margin note could consist of text that is 100% generated by AI, or could involve more human-generated text.
@@
Third, referring back to specific responses requires additional work. For example, suppose a user prompted a chat-based LLM interface to get an explanation for an unfamiliar concept in a document. The explanation may be remembered in the short term, but if the user revisits the document much later, they must scroll through a lengthy and poorly organized chat history to find it, which may be especially difficult for readers with a lower working memory capacity to do (ToScrollOrNot2009). Linking responses directly to the relevant text avoids this issue. Some systems, like Adobe Acrobat Reader, support clicking on chat responses to highlight relevant parts of the document, however, this still requires scrolling through chat history. Alternatively, the user could re-issue the prompt, but this wastes time and resources. Of course, users could copy responses they wish to save and paste them in comments (kreijkes2025effects), but this requires additional copy and paste steps that would not be needed with an AI margin note.
@@ 3.2. Selection Automation
AI margin notes are associated with specific text, which defines the context for the prompt and a location to display the note. Specifying text for the note could be done automatically by the AI assistant or manually by the user (Figure 3). For example, consider summarization, a common task readers do in conventional margin notes (jackson2001marginalia; porter2004beyond) and with LLM systems (melin2024ai; kreijkes2025effects). When using LLMs in document reader software, readers can generate a summary of the entire document automatically by pressing a single button (AdobeAcrobatAI), or they can manually specify which parts of the document need to be summarized by copying and pasting text into the chat. When the reader just presses a button, the LLM decides which information is important and worth including in the summary, but when the reader copies text to summarize, they are deciding which parts of the document are most important. Prior work on text highlighting suggests that this decision process could improve learning outcomes and reading comprehension, but it could also increase mental effort (yue2015highlighting; joshi2024constrained).
Automating the selection of text to create AI margin notes may also impact psychological ownership, feelings of the learning experiences belonging to them (Pierce2001PsychologicalOwnership). Prior work shows that fostering psychological ownership, which can be achieved by giving learners more control over their learning (Pierce2001PsychologicalOwnership; Pierce2003PsychologicalOwnership), can encourage more active learning (griffith1990cooperative) and motivates learning (soslau2007urban). Automatically generating AI margin notes reduces effort, but potentially at the risk of lowering comprehension and psychological ownership.
@@ 3.3. Human and AI Involvement
Currently, commenting in document reader software is intended for text that is entirely written by a human. Specifically, the user must formulate their own ideas and type into a text box. At the other extreme, pressing a “Summarize” button in a document reader (AdobeAcrobatAI) produces text that is entirely written by the LLM without any guidance from the human. However, there are techniques that require involvement from both the human and the LLM. For example, a chat-based interface requires the user to provide a specific instruction to the LLM, such as “summarize the text for someone in high school.” Here, the user forms goals and sub-tasks (MetacognitiveDemandsPrompting2024) to exert some control over the output (ConstrainedGPT), and the LLM produces it. These roles can be reversed too, for example, the LLM could ‘prompt’ the user to respond to practice reading comprehension questions (wang2024chatprcs; liu2023student).
AI margin notes can also vary in how much human and AI involvement they require to produce the final comment text (Figure 4). Techniques that require more human involvement likely require more cognitive engagement than those with more AI involvement. This may be beneficial for reading comprehension (bjork1994memory; bjork2011making; bjork1994institutional; tyler1979cognitive), may discourage readers from taking verbatim notes (bretzing1979processing), and may even improve feelings of psychological ownership (ConstrainedGPT; joshi2025interactiontechniquesencouragelonger). However, too much cognitive engagement can frustrate and discourage learners (grund2024learning), which may be mitigated by increasing AI involvement.